Graziano Corporation (GC) is considering a project to purchase new equipment. The equipment would be depreciated by the straight-line method over its 3-year life and would have a zero-salvage value. The project requires an investment of $6,000 today on net working capital. Revenues and other operating costs are expected to be constant over the project's 3-year life. However, this project would compete with other company’s products and would reduce its pre-tax annual cash flows of $5,000 per year. The investment on the project costs $76,000 today plus $4,000 installation cost. The project generates sales revenues of $67,500 each year for 3 years. The annual operating costs of the project is $25,000, excluding depreciation. GC has a weighted average cost of capital 11.00% and its tax rate is 35 percent. The company can sell the equipment at the end of the third year to generate $10,000 after-tax cash flow. What is the project's MIRR and NPV Graziano Corporation (GC) is considering a project to purchase new equipment. The equipment would be depreciated by the straight-line method over its 3-year life and would have a zero-salvage value. The project requires an investment of $6,000 today on net working capital. Revenues and other operating costs are expected to be constant over the project's 3-year life. However, this project would compete with other company’s products and would reduce its pre-tax annual cash flows of $5,000 per year. The investment on the project costs $76,000 today plus $4,000 installation cost. The project generates sales revenues of $67,500 each year for 3 years. The annual operating costs of the project is $25,000, excluding depreciation. GC has a weighted average cost of capital 11.00% and its tax rate is 35 percent. The company can sell the equipment at the end of the third year to generate $10,000 after-tax cash flow. What is the project's MIRR and NPV
Answer #1
1] Cost of investment+installation = 76000+4000 = $              80,000.00
Increase in NWC $                6,000.00
Initial outlay $              86,000.00
2] Annual revenue from the project $              67,500.00
-Annual operating costs other than depreciation $              25,000.00
-Depreciation = 80000/3 = $              26,666.67
-Reduction in other products' pretax cash flow $                5,000.00
-Incremental NOI $              10,833.33
-Tax at 35% $                3,791.67
=Incremental NOPAT $                7,041.67
+Depreciation $              26,666.67
=Incremenal OCF $              33,708.33
3] After tax cash outflow from sale of equipment $              10,000.00
Recovery of NWC $                6,000.00
Terminal non operating cash flows $              16,000.00
4] PV of incremental OCF = 33708.33*(1.11^3-1)/(0.11*1.11^3) = $              82,373.54
PV of terminal non operating cash flows = 16000/1.11^3 = $              11,699.06
Sum of PV of cash inflows $              94,072.60
Less: Initial investment $              86,000.00
NPV $                8,072.60
5] MIRR assumes that intervening cash flows are invested at WACC. Hence, the
FV of the cash inflows, namely the annual incremental OCF and the terminal
non-operating cash flows are calculated. It is done below:
FV of incremental OCF = 33708.33*(1.11^3-1)/(0.11) = $          1,12,656.61
FV of terminal non operating cash flows $              16,000.00
Sum of FV of cash inflows $          1,28,656.61
Now we have only two cash flows for the project.
1] The sum of FVs of the inflows which, is equal to $128,656.61 at t3, and
2] The initial cash outflow of $86,000 at t0.
Now MIRR is the discount rate that equals the two.
That is 86000 = 128656.61/(1+MIRR)^3
Hence,
MIRR = (128656.61/86000)^(1/3)-1 = 14.37%